OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, –3) and has a slope of 1/2?

OpenStudy (anonymous):

just substitute these values into point slope equations and then solve for standard form

OpenStudy (anonymous):

Equation is given by: \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

Here: \((x_1, y_1) = (-2, -3)\) and m = 1/2 Just plug in and solve..

OpenStudy (anonymous):

y + 3 = 1/2(x + 3)

OpenStudy (anonymous):

y + 3 = 1/2x + 3/2

OpenStudy (anonymous):

y = 1/2x + 9/2

OpenStudy (anonymous):

how do i put that in standard form?

OpenStudy (anonymous):

x1 is -2 here..

OpenStudy (anonymous):

oh i see where i messed up!

OpenStudy (anonymous):

\[y + 3 = \frac{1}{2}(x+2)\] Now here multiply by 2 both the sides and tell me what did you get??

OpenStudy (anonymous):

2y + 6 = x + 1

OpenStudy (anonymous):

There is no need to go fast.. Do it slowly.. Is your RHS right ??

OpenStudy (anonymous):

rhs?

OpenStudy (anonymous):

Right hand side..

OpenStudy (anonymous):

i think so.

OpenStudy (anonymous):

How did you get x + 1 ??

OpenStudy (anonymous):

oh it's x + 2

OpenStudy (anonymous):

Standard form says that x and y should be on one side and constants on other side.. can you do that ??

OpenStudy (anonymous):

\(2y + 6 = x + 2\) Subtract 2y both the sides here..

OpenStudy (anonymous):

6= -2y + x + 2

OpenStudy (anonymous):

4 = -2y + x

OpenStudy (anonymous):

Yes.. Now subtract 2 both the sides..

OpenStudy (anonymous):

Now rearrange it to look better..

OpenStudy (anonymous):

-2y + x = 4

OpenStudy (anonymous):

i mean x – 2y = 4

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

thanks dog

OpenStudy (anonymous):

Welcome..